OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –2 7 0 6 Line 2 x y 3 0 0 –3 A.0 B.exactly 1 C.exactly 2 D.infinitely many The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –4 3 –2 4 Line 2 x y 0 8 2 6 A.0 B.exactly 1 C.exactly 2 D.infinitely many plz help. will fan and medal

OpenStudy (bohotness):

okay. so, first you need the slope of both lines. Slope equation is (y2-y1)/(x2-x1) (Line 1) 2--2/1--1 4/2 2 (Line 2)

OpenStudy (bohotness):

(1-10)/(4--2) -9/6 -3/2

OpenStudy (calculusfunctions):

I). Infinitely many solutions if both lines have the same slope and the same intercept(s). II). No solution if they have the same slope but different intercepts. III). One solution if they have different slopes.

OpenStudy (bohotness):

Then you can put this into slope-intercept form. y=mx+b (Line 1) -2 = 2(-1) + b -2 = -2 +b -2+ 2 = b b = 0 so line 1's equation is y = 2x

OpenStudy (calculusfunctions):

@bohotness you're supposed to teach! Not give out the answers!

OpenStudy (bohotness):

iam not giving answer

OpenStudy (calculusfunctions):

You just did the steps above. Hence if you provide full or part of the solution, then you have given the answer!